Solve the following system of linear equations. (If the system has no solution, enter NONE for each answer. If the system is dependent, enter INFINITELY MANY for each answer.)
6x-2y=7
x-y=2
x = ?
y = ?
pls show me the steps to get the answer

Answer by mananth(16946)

6x-2y=7 .............1
x-y=2 .............2

multiply (1)by 1
Multiply (2) by -2
6x-2y=7
-2x+2y= -4

Add the two equations

4x=3

/4

x= 3/4
plug value of x in (1)

6x-2y=7
4.5-2y =7
-2y=7
-2y=2.5
y=-1 1/4